# A galactic outflow traced by its extended Mg II emission out to a $\sim 30$ kpc radius in the Hubble Ultra Deep Field with MUSE

### v1

@@ 7.2 Spectra extracted from annular apertures
Beyond comparing the Mg II emission pseudo-narrow band images for the data and model cubes, it is also relevant to compare their respective spectral energy distributions. To do this, we extract and compare the spectra from the same ring-like apertures defined in Sec. 4.2 and shown in Fig. 3, for the data cube and the best-fitting model cube.
Figure 11: Rest-frame spectra extracted from five ring-like apertures as well as from the full modeled field of view, for the data cube (solid lines) and the best-fitting model cube (dashed lines). The labels in each panel indicate the inner and outer radii of each ring-like aperture (except the innermost, circular aperture, which is defined by a single radius). In the frame of reference of the central galaxy, the apertures represent physical sizes of $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc. The nearly horizontal dotted lines show the continuum level for each aperture. The vertical dotted lines indicate the rest-frame wavelength of the Mg II doublet. The smaller panels show the residuals of the model, normalized by the median standard deviation of the spectra extracted from each aperture. That is, a residual of $1$ indicates one sigma difference between the data and the model. Our outflow modeling scheme reproduces the Mg II emission and absorption profiles relatively well. The top-left panel shows in gray colors the absorption- and emission-dominated bands (light gray and dark gray, respectively), defined by the wavelengths where the full FoV model intersects the continuum. We use these bands to produce emission- and absorption-narrow band images that we analyze in Sec. 7.3
Figure 11 shows the spectra extracted from these apertures, as well as from the full modeled field of view, for the data (solid lines) and the model (dashed lines) cubes. It is clear that there is relatively good agreement between both cubes, with an inner spectrum dominated by strong absorption features and little emission, and a progressively more emission-dominated spectrum towards the outer apertures. The model successfully reproduces the observed line profiles, especially in the inner apertures. On the other hand, there are some differences between the data and the model that become more obvious in the outer apertures. For instance, the model does not reach the level of emission present in the data cube in the outer apertures. The observed emission lines in the outer regions are slightly but systematically stronger than in the model. Furthermore, the emission in the outer apertures of the data cube is slightly blue-shifted with respect to the inner apertures, and shows in general a more complex kinematic structure. Some of these differences are discussed more in detail in Sec. 8.2. Recall that the radial binning of the data and model cubes is only for visualization and comparison purposes. The model is fitted to the data cube, not to aperture-integrated spectra.
@@ 7.3 Radial profiles
In the top-left panel of Fig. 11 we show in gray the absorption- and emission-dominated bands, defined by the wavelengths where the continuum intersects the modeled P-Cyngi profile, for the full FoV spectrum. To compare the spatial distribution of Mg II emission between the model and the data, we have computed the absorption- and emission-centered narrow band images, summing the wavelength channels within these absorption- and emission-dominated bands, for the continuum-subtracted data and model cubes.
@@
Figure 12: Radial profiles of the Mg II surface brightness measured for the absorption (blue) and emission (red) dominated bands, defined in the top-left panel of Fig. 11, as well as for the full wavelength range (i.e., total absorption plus emission), for the data cube (solid lines) and best-fitting model cube (dashed lines). The brown dashed line indicates the zero level (i.e., continuum surface brightness). The model reproduces several key features of the radial profiles, such as a similar depth in the absorption in the central region, and a decrease in the emission profile towards outer radii generally consistent within the error bars, with a peak at around $\sim 5$ kpc from the central galaxy.
@@ 8.1.1 Single scattering approximation
As explained in Sec. 5.1, if the velocity field is steep enough, the Mg II photons are only able to interact with the ions in a specific discrete spherical shell, and they can thus easily escape immediately after being reemitted. On the other hand, if the velocity field is closer to flat (i.e., $\gamma\approx 0$), the intrinsic velocity distribution of ions within a given spherical shell (or also the presence of turbulence) would allow the Mg II photons to interact with ions within a broader layer, thus, making the reabsorption of photons more likely. However, since Mg II does not present an alternative fluorescent channel, these reabsorbed photons would eventually be reemitted again through the same transition, although possibly at a different wavelength than the originally reemitted photon. This could potentially modify the line shape of the emission component of the P-Cygni profile.
While implementing a more realistic treatment to model multiple scattering within the outflow is beyond the scope of this paper, as it would require the implementation of full radiative transfer numerical simulations (see, e.g., Prochaska et al., 2011), our results suggest a velocity field with a reasonably well-constrained steepness $\gamma\approx 0.7$, supporting the idea that the single scattering approximation is a justified choice for this case (our approach does not prevent at any point a flatter velocity field, therefore, the steep gradient obtained is not a consequence of our assumptions).
@@ 8.1.2 Complexity of CGM physics
Furthermore, a combination of observations and simulations suggests that outflowing material from the CGM presents a spatial distribution consistent with clumps of gas varying in strength on physical scales ¡ 2 kpc (see, e.g., Augustin et al., 2021). Similarly, Afruni et al. (2023) use a sample of Mg II $\lambda\lambda$2796,2803 absorbers in the MUSE spectra of giant gravitational arcs to infer a coherence scale of the gas in the CGM in the range of $1.4-7.8$ kpc, and Nelson et al. (2020) report similar conclusions using purely cosmological magnetohydrodynamical simulations. This contrasts with the smoothly varying gas density distribution assumed in our model. Hence, reproducing small-scale features of the CGM is beyond the scope of our model, although this becomes less important when we compare spatially-averaged quantities (e.g., aperture-integrated spectra), such as in Sec. 7.3.
@@ 8.2 Comparing data and model cubes
  • The model does not accurately reproduce the inner radial profile of the emission-dominated band image (see the right panel of Fig. 12). The model exhibits excess emission that peaks at about $\sim 5$ kpc from the central source. Although this difference could, in principle, be caused by noise, since the inner radial bins are more strongly affected by noise due to the lower number statistics (i.e., fewer pixels), it is more likely related to the model not reproducing the observed relative strength of the two Mg II transitions, putting a relative excess of flux in the red line of the doublet as a result (see also the top right panel of Fig. 11). Although by construction, we use half of the optical depth $\tau_{0}$ for the Mg II $\lambda$2803 line, this hardly makes any difference at such high overall optical depth (due to the factor $e^{-\tau}$). Thus, a stronger Mg II $\lambda$2796 emission in the data (compared to the Mg II $\lambda$2803 line) could hint at unaccounted-for nebular emission. The fact that in the outer apertures, where the continuum contribution is virtually zero, the ratio between the two lines is closer to one provides further evidence that this is the case. This would be in agreement with the findings from Seon (2023) for an expanding spherical outflow, where the authors use radiative transfer simulations to explore radial variations in the Mg II $\lambda$2796-to-$\lambda$2803 ratio induced by resonant scattering. As for why is this contribution not fully captured by our model, an explanation could be that it is of second order, compared to the continuum scattered emission, and harder to constrain given our flux uncertainties, although our best-fitting model already accounts for some nebular contribution. In addition, differences in orientation and occultation of light on disk scales could also contribute to enhancing discrepancies between the data and the model in the inner part of the radial profiles.
@@
  • The data cube presents stronger emission in the outer parts ($r>9$ kpc), compared to the model cube. A discrepancy in the emission towards larger radii is not unexpected, since an incomplete description of the gas kinematics (as discussed above) also yields differences in total flux. This is because the gas density (and, thus, optical depth) depends primarily on the velocity field assumed (through Eq. 2).
  • The observed Mg II halo does not present a smooth spatial distribution. This is evident in Fig. 3 for the outer parts of the halo, and in Fig. 13 for the inner parts. Instead, it exhibits a more irregular distribution, consistent with previous findings from observations and simulations (see discussion in Sec. 8.1.2). However, our model, as described in Sec. 5, is not able to reproduce such a configuration due to the underlying spherical symmetry assumed in Eq. 2.
@@ 8.3 Strong Mg II absorption at 30 kpc from the galaxy
As mentioned in Sec. 3.1, the spectrum of the background galaxy BG1 ($z=1.221$, $m_{\mathrm{AB,F775W}}=24.5$, projected impact parameter to #884 of 30 kpc) shows strong Mg II doublet absorption at the same redshift as #884.
@@
This comes with the caveat that the galaxy BG1 is beyond the modeled radius, so this implies an extrapolation of the model. In previous sections we have established a number of differences between the data and model, that are mostly relevant in the outer apertures. Hence, BG1 provides an independent sightline to test our model and find out what the discrepancies are between the data and the model towards even larger impact parameters, where the emission of Mg II is more difficult to probe due to contamination from nearby galaxies (such as BG1) and the overall lower S/N. While our model was adjusted to fit the data cube at impact parameters $<26$ kpc, it can, in principle, be used to infer the distribution of the outflowing gas toward larger distances, and evaluate if a simple outflow model is still a valid approximation to describe the outer CGM.
@@
Thus, we evaluate our model at the BG1 sightline to check if the observed absorption strength is consistent with our predicted value. Figure 14 shows the Mg II doublet absorption at the redshift of #884 in the spectrum of the galaxy BG1. The absorption lines are nearly completely saturated, and slightly redshifted compared to the absorption present in the #884 spectrum, by about $65$ km s-1. The absorption in the spectrum of BG1 is also considerably broader than that present in the spectrum of #884. This is because in the case of BG1, both sides of the biconical outflow absorb flux, instead of only the blue-shifted half of the outflow, as for #884. The figure shows that although the absorption strength predicted by our model (red) is not negligible (more than $\sim 70\%$ of the continuum), it is significantly weaker than the observed absorption. In Sec. 11 we show that our model underestimates the total emission (hence, gas density) in the outer apertures. Therefore, it is not surprising that it also underestimates the absorption strength towards BG1 at an impact parameter of $\sim 30$ kpc.
@@
Additionally, the modeled absorption is considerably broader than the observed. This hints possibly at an overestimation of the maximum radius by the model. Indeed, a maximum radius of $40$ kpc (which would correspond to a $V_{\mathrm{max}}$ of $\sim 330$ km s-1, keeping other parameters unchanged) produces absorption features that match the width of the absorption in the spectrum of BG1 much better. This value is, however, significantly lower than the $V_{\mathrm{max}}$ derived from the inner regions of the outflow (although $40$ kpc is indeed consistent within the error bars with our inferred $R_{\mathrm{max}}=70\pm 40$ kpc), suggesting again that our simple prescription to model the velocity of the gas is not able to describe the more complex kinematics/morphology of the outer regions of the CGM of #884.
@@
Finally, the biconical geometry of the modeled outflow, that produces the asymmetric line profiles, does not provide an accurate description of the absorption features at this large impact parameter. Qualitatively, the absorption features are closer to symmetric, hinting to a spherical configuration of the CGM in the outer region (or at least, a more continuous distribution of gas along this discrete sightline). An alternative could be that this absorption is probing a cold and spatially coherent inflow of gas (this idea would also be consistent with the fact that BG1 is located nearly perpendicular to the major axis of the outflow in Fig. 3, see, e.g., Weng et al., 2023)
@@
Figure 14: Mg II doublet absorption at the redshift of #884 in the spectrum of the galaxy BG1 (black). The horizontal dotted line shows the continuum level in the spectrum of BG1. The vertical dotted lines show the rest-frame wavelengths of the Mg II doublet. The red dashed line shows the absorption line predicted by extrapolating our best-fitting model to the location of the BG1 sightline. The blue dashed line shows the full FoV spectrum of #884, with the continuum level renormalized to match that of BG1. The absorption in the spectrum of BG1 is considerably broader than that present in the spectrum of #884. This is because in the case of BG1, both sides of the spherical outflow absorb flux, instead of only the blue-shifted half of the outflow. The predicted absorption is inconsistent with the observed profiles, implying that our outflow model does not describe the gas distribution properly along this sightline.
@@ 8.4 Mass outflow rate and mass loading factor
Previous works have aimed at inferring the amount of outflowing gas by estimating the radius at which the optical depth becomes $\sim 1$, and the velocity gradient of the outflow from the width of the observed spectral lines (e.g., Martin et al., 2013; Guo et al., 2023a). In this analysis, our outflow model provides a self-consistent mass outflow rate that results from the combination of the velocity law and density/optical depth profile described in Sec. 5.1. Thus, at any point in the 3-D space, we can infer the density and the velocity of the gas from our analytic framework. Here, we present and discuss the implications of the constraints on the mass outflow rate inferred from our outflow model.
The density of the outflowing gas is directly proportional to the inferred optical depth through Eq. 6. Replacing the constants, we have:
$\tau_{0}=4.57\times 10^{-7}\mathrm{cm}^{3}\mathrm{s}^{-1}n_{0}\frac{R_{0}}{v_{0}}$ (25)
which leads to $n_{0}=5.76\times 10^{-6}$ cm-3, using the best-fitting model values of $\tau_{0}$, $R_{0}$, and $v_{0}$. However, in this context, $n_{0}$ corresponds to the number density of ionized Mg ($n_{\mathrm{Mg^{+}},0}$) at radius $R_{0}$. To compute the corresponding hydrogen density of the gas ($n_{\mathrm{H},0}$), it is necessary to correct by metallicity $\eta(\mathrm{Mg})$, dust depletion $d(\mathrm{Mg})$ and ionization fraction $\chi(\mathrm{Mg^{+}})$ (see, e.g., Martin et al., 2013). Specifically, $n_{\mathrm{H},0}$ is given by:
$n_{\mathrm{H},0}=\frac{n_{\mathrm{Mg^{+}},0}}{\eta(\mathrm{Mg})\,\chi(\mathrm{Mg^{+}})\,d(\mathrm{Mg})}$ (26)
@@
We use then the relation between $\chi(\mathrm{Mg^{+}})$ and $U$ reported by Martin et al. (2013), determined using Cloudy (Ferland et al., 2013), to compute the corresponding ionization fraction:
$\chi(\mathrm{Mg^{+}})=4\times 10^{-4}U^{-0.94}$ (27)
@@
$\dot{M}(r)=\mu\,m_{\mathrm{H}}\,n_{\mathrm{H}}(r)\,A(r)\,v(r)$ (28)
with $\mu\,m_{\mathrm{H}}$ equal to the average mass per hydrogen atom. We adopt a value of $\mu=1.4$, which is commonly used in this context (see, e.g., Rupke et al., 2002; Weiner et al., 2009; Martin et al., 2012; Burchett et al., 2021). Since we have imposed mass conservation, the expression above does not depend on radius and can be written as:
$\dot{M}=4\,\pi\,\mu\,m_{\mathrm{H}}\,n_{\mathrm{H},0}\,C_{\Omega}\,R_{0}^{2}\,v_{0}$ (29)
where $C_{\Omega}=1-cos(\mathrm{O.A.})$ corresponds to the angular covering fraction of the bicone. Evaluating the equation yields a mass outflow rate of:
$\dot{M}\approx 12\pm 7\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1}$ (30)
@@
Overall, the measured outflow rate and mass loading factor could be up to a factor $\sim 6$ lower, if we assume a gas-phase metallicity given by the relation reported by Zahid et al. (2014), and boosted according to the findings from Chisholm et al. (2018) (with the caveat that a different metallicity calibration could potentially shift the relation from Zahid et al. (2014)). On the contrary, it could also be up to a factor $\sim 8$ higher, using the dust depletion reported by Savage & Sembach (1996). Despite that we are not accounting for additional systematic variations induced by our ionization correction choice, the metallicity and dust depletion choices come with significant systematic uncertainties that must be taken into account in order to properly interpret our results and compare them with the results from different works.
@@
In the following, we proceed similarly to Burchett et al. (2021) to investigate whether this outflowing material is able to escape the galactic potential or will fall again into the galaxy. We use the stellar-to-halo mass relation from Moster et al. (2013) to calculate a total halo mass of $\log M_{\mathrm{h}}/\mathrm{M}_{\odot}=11.9$. For halo of this mass with a Navarro–Frenk–White profile, and assuming a concentration parameter given by the relation from Ragagnin et al. (2019), we calculate an escape velocity of $\sim 250$ km s-1 at the maximum radius reached by the outflow of $\sim 70$ kpc. This is significantly lower than the terminal velocity of the outflow $v_{\mathrm{max}}=450$ km s-1, meaning that the outflowing material will likely escape the gravitational potential of the halo and enrich the intergalactic medium, rather than falling again into #884. However, as shown in previous sections, a radius of $70$ kpc is well beyond the radius where our outflow model describes the data faithfully. If we compare the escape velocity with the outflow velocity at a smaller radius, where our outflow model still performs well, we find that, for instance, at $25$ kpc, the escape velocity is still higher (although comparable, within the error bars) than the outflow velocity ($\sim 250$ and $\sim 220\pm 70$ km s-1, respectively). Specifically, given our set of best-fitting parameters, we find that the outflow starts exceeding the escape velocity of the galactic potential at $\sim 30$ kpc. Thus, whether the outflowing gas escapes the galactic potential or not depends on the radii up to which our outflow model provides an accurate representation of the data. Our analysis suggests that this holds for impact parameters of $\sim 20-25$ kpc. However, this cannot be translated directly into radii due to the additional line-of-sight component of the distance, and the face-on nature of the outflow. Moreover, we find evidence of high-velocity gas ($\sim 400$ km s-1) in the down-the-barrel spectrum of the outflow. Altogether, we find it more likely that our outflow model still performs well by the point where the outflow velocity overcomes the escape velocity of the galaxy halo and, thus, that the outflowing gas will be able to escape the galaxy.

### v3

@@ 7.2 Spectra extracted from annular apertures
Beyond comparing the Mg II emission pseudo-narrowband images for the data and model cubes, it is also relevant to compare their respective spectral energy distributions. To do this, we extract and compare the spectra from the same ring-like apertures defined in Sect. 4.2 and shown in Fig. 3, for the data cube and the best-fitting model cube.
Figure 11: Rest-frame spectra extracted from five ring-like apertures as well as from the full modeled FoV for the data cube (solid lines) and the best-fitting model cube (dashed lines). The labels in each panel indicate the inner and outer radii of each ring-like aperture (except the innermost, circular aperture, which is defined by a single radius). In the frame of reference of the central galaxy, the apertures represent physical sizes of $<3$ kpc, $3-6$ kpc, $6-9$ kpc, $9-15$ kpc, and $15-26$ kpc. The nearly horizontal dotted lines show the continuum level for each aperture. The vertical dotted lines indicate the rest-frame wavelength of the Mg II doublet. The smaller panels show the residuals of the model, normalized by the median standard deviation of the spectra extracted from each aperture. That is, a residual of $1$ indicates one sigma difference between the data and the model. Our outflow modeling scheme reproduces the Mg II emission and absorption profiles relatively well. The top-left panel shows in gray colors the absorption- and emission-dominated bands (light gray and dark gray, respectively), defined by the wavelengths where the full FoV model intersects the continuum. We use these bands to produce emission- and absorption-narrowband images that we analyze in Sect. 7.3
Figure 11 shows the spectra extracted from these apertures, as well as from the full modeled FoV, for the data (solid lines) and the model (dashed lines) cubes. It is clear that there is relatively good agreement between both cubes, with an inner spectrum dominated by strong absorption features and little emission, and a progressively more emission-dominated spectrum toward the outer apertures. The model successfully reproduces the observed line profiles, especially in the inner apertures. On the other hand, there are some differences between the data and the model that become more obvious in the outer apertures. For instance, the model does not reach the level of emission present in the data cube in the outer apertures. The observed emission lines in the outer regions are slightly but systematically stronger than in the model. Furthermore, the emission in the outer apertures of the data cube is slightly blue-shifted with respect to the inner apertures, and shows in general a more complex kinematic structure. Some of these differences are discussed more in detail in Sect. 8.2. Recall that the radial binning of the data and model cubes is only for visualization and comparison purposes. The model is fitted to the data cube, not to aperture-integrated spectra.
@@ 7.3 Radial profiles
In the top-left panel of Fig. 11 we show in gray the absorption- and emission-dominated bands, defined by the wavelengths where the continuum intersects the modeled P-Cyngi profile, for the full FoV spectrum. To compare the spatial distribution of Mg II emission between the model and the data, we have computed the absorption- and emission-centered narrow-band images, summing the wavelength channels within these absorption- and emission-dominated bands, for the continuum-subtracted data and model cubes.
@@
Figure 12: Radial profiles of the Mg II surface brightness measured for the absorption (blue) and emission (red) dominated bands, defined in the top-left panel of Fig. 11, as well as for the full wavelength range (i.e., total absorption plus emission), for the data cube (solid lines) and best-fitting model cube (dashed lines). The brown dashed line indicates the zero level (i.e., continuum surface brightness). The model reproduces several key features of the radial profiles, such as a similar depth in the absorption in the central region, and a decrease in the emission profile toward outer radii generally consistent within the error bars, with a peak at around $\sim 5$ kpc from the central galaxy.
@@ 8.1.1 Single scattering approximation
As explained in Sect. 5.1, if the velocity field is steep enough, the Mg II photons are only able to interact with the ions in a specific discrete spherical shell, and they can thus easily escape immediately after being reemitted. On the other hand, if the velocity field is closer to flat (i.e., $\gamma\approx 0$), the intrinsic velocity distribution of ions within a given spherical shell (or also the presence of turbulence) would allow the Mg II photons to interact with ions within a broader layer, thus, making the reabsorption of photons more likely. However, since Mg II does not present an alternative fluorescent channel, these reabsorbed photons would eventually be reemitted again through the same transition, although possibly at a different wavelength than the originally reemitted photon. This could potentially modify the line shape of the emission component of the P-Cygni profile.
While implementing a more realistic treatment to model multiple scattering within the outflow is beyond the scope of this paper, as it would require the implementation of full radiative transfer numerical simulations (see, e.g., Prochaska et al., 2011), our results suggest a velocity field with a reasonably well-constrained steepness $\gamma\approx 0.7$. This supports the idea that the single scattering approximation is a justified choice for this case (our approach does not prevent at any point a flatter velocity field, and therefore the steep gradient obtained is not a consequence of our assumptions).
@@ 8.1.2 Complexity of CGM physics
Furthermore, a combination of observations and simulations suggests that outflowing material from the CGM presents a spatial distribution consistent with clumps of gas varying in strength on physical scales ¡ 2 kpc (see, e.g., Augustin et al., 2021). Similarly, Afruni et al. (2023) use a sample of Mg II $\lambda\lambda$2796,2803 absorbers in the MUSE spectra of giant gravitational arcs to infer a coherence scale of the gas in the CGM in the range of $1.4-7.8$ kpc, and Nelson et al. (2020) report similar conclusions using purely cosmological magnetohydrodynamical simulations. This contrasts with the smoothly varying gas density distribution assumed in our model. Hence, reproducing small-scale features of the CGM is beyond the scope of our model, although this becomes less important when we compare spatially averaged quantities (e.g., aperture-integrated spectra), such as in Sect. 7.3.
@@ 8.2 Comparing data and model cubes
The model does not accurately reproduce the inner radial profile of the emission-dominated band image (see the right panel of Fig. 12). The model exhibits excess emission that peaks at about $\sim 5$ kpc from the central source. Although this difference could, in principle, be caused by noise, since the inner radial bins are more strongly affected by noise due to the lower number statistics (i.e., fewer pixels), it is more likely related to the model not reproducing the observed relative strength of the two Mg II transitions, putting a relative excess of flux in the red line of the doublet as a result (see also the top right panel of Fig. 11). Although by construction, we use half of the optical depth $\tau_{0}$ for the Mg II $\lambda$2803 line, this hardly makes any difference at such high overall optical depth (due to the factor $e^{-\tau}$). Thus, a stronger Mg II $\lambda$2796 emission in the data (compared to the Mg II $\lambda$2803 line) could hint at unaccounted-for nebular emission. The fact that in the outer apertures, where the continuum contribution is virtually zero, the ratio between the two lines is closer to one provides further evidence that this is the case. This would be in agreement with the findings from Seon (2024) for an expanding spherical outflow, where the authors use radiative transfer simulations to explore radial variations in the Mg II $\lambda$2796-to-$\lambda$2803 ratio induced by resonant scattering. As for why this contribution is not fully captured by our model, an explanation could be that it is of second order, compared to the continuum scattered emission, and harder to constrain given our flux uncertainties, although our best-fitting model already accounts for some nebular contribution. In addition, differences in orientation and occultation of light on disk scales could also contribute to enhancing discrepancies between the data and the model in the inner part of the radial profiles.
@@
The data cube presents stronger emission in the outer parts ($r>9$ kpc), compared to the model cube. A discrepancy in the emission toward larger radii is not unexpected, since an incomplete description of the gas kinematics (as discussed above) also yields differences in total flux. This is because the gas density (and, thus, optical depth) depends primarily on the velocity field assumed (through Eq. 2).
The observed Mg II halo does not present a smooth spatial distribution. This is evident in Fig. 3 for the outer parts of the halo, and in Fig. 13 for the inner parts. Instead, it exhibits a more irregular distribution, consistent with previous findings from observations and simulations (see discussion in Sect. 8.1.2). However, our model, as described in Sect. 5, is not able to reproduce such a configuration due to the underlying spherical symmetry assumed in Eq. 2.
@@ 8.3 Strong Mg II absorption at 30 kpc from the galaxy
As mentioned in Sect. 3.1, the spectrum of the background galaxy BG1 ($z=1.221$, $m_{\mathrm{AB,F775W}}=24.5$, projected impact parameter for #884 of 30 kpc) shows strong Mg II doublet absorption at the same redshift as #884.
@@
This comes with the caveat that the galaxy BG1 is beyond the modeled radius, so this implies an extrapolation of the model. In previous sections we have established a number of differences between the data and model, that are mostly relevant in the outer apertures. Hence, BG1 provides an independent sight line to test our model and find out what the discrepancies are between the data and the model toward even larger impact parameters, where the emission of Mg II is more difficult to probe due to contamination from nearby galaxies (such as BG1) and the overall lower S/N. While our model was adjusted to fit the data cube at impact parameters $<26$ kpc, it can, in principle, be used to infer the distribution of the outflowing gas toward larger distances, and evaluate if a simple outflow model is still a valid approximation to describe the outer CGM.
@@
Thus, we evaluate our model at the BG1 sight line to check if the observed absorption strength is consistent with our predicted value. Figure 14 shows the Mg II doublet absorption at the redshift of #884 in the spectrum of the galaxy BG1. The absorption lines are nearly completely saturated, and slightly redshifted compared to the absorption present in the #884 spectrum, by about $65$ km s-1. The absorption in the spectrum of BG1 is also considerably broader than that present in the spectrum of #884. This is because in the case of BG1, both sides of the biconical outflow absorb flux, instead of only the blue-shifted half of the outflow, as for #884. The figure shows that although the absorption strength predicted by our model (red) is not negligible (more than $\sim 70\%$ of the continuum), it is significantly weaker than the observed absorption. In Sect. 11, we show that our model underestimates the total emission (hence, gas density) in the outer apertures. Therefore, it is not surprising that it also underestimates the absorption strength toward BG1 at an impact parameter of $\sim 30$ kpc.
@@
Additionally, the modeled absorption is considerably broader than the observed. This hints possibly at an overestimation of the maximum radius by the model. Indeed, a maximum radius of $40$ kpc (which would correspond to a $V_{\mathrm{max}}$ of $\sim 330$ km s-1, keeping other parameters unchanged) produces absorption features that match the width of the absorption in the spectrum of BG1 much better. This value is, however, significantly lower than the $V_{\mathrm{max}}$ derived from the inner regions of the outflow (although $40$ kpc is indeed consistent within the error bars with our inferred $R_{\mathrm{max}}=70\pm 40$ kpc), suggesting again that our simple prescription to model the velocity of the gas is not able to describe the more complex kinematics and/or morphology of the outer regions of the CGM of #884.
@@
Finally, the biconical geometry of the modeled outflow, that produces the asymmetric line profiles, does not provide an accurate description of the absorption features at this large impact parameter. Qualitatively, the absorption features are closer to symmetric, hinting to a spherical configuration of the CGM in the outer region (or at least, a more continuous distribution of gas along this discrete sight line). An alternative could be that this absorption is probing a cold and spatially coherent inflow of gas (this idea would also be consistent with the fact that BG1 is located nearly perpendicular to the major axis of the outflow in Fig. 3; see, e.g., Weng et al., 2024)
@@
Figure 14: Doublet absorption of Mg II at the redshift of #884 in the spectrum of the galaxy BG1 (black). The horizontal dotted line shows the continuum level in the spectrum of BG1. The vertical dotted lines show the rest-frame wavelengths of the Mg II doublet. The red dashed line shows the absorption line predicted by extrapolating our best-fitting model to the location of the BG1 sight line. The blue dashed line shows the full FoV spectrum of #884, with the continuum level renormalized to match that of BG1. The absorption in the spectrum of BG1 is considerably broader than that present in the spectrum of #884. This is because in the case of BG1, both sides of the spherical outflow absorb flux, instead of only the blue-shifted half of the outflow. The predicted absorption is inconsistent with the observed profiles, implying that our outflow model does not describe the gas distribution properly along this sight line.
@@ 8.4 Mass outflow rate and mass loading factor
Previous works have aimed at inferring the amount of outflowing gas by estimating the radius at which the optical depth becomes $\sim 1$, and the velocity gradient of the outflow from the width of the observed spectral lines (e.g., Martin et al., 2013; Guo et al., 2023a). In this analysis, our outflow model provides a self-consistent mass outflow rate that results from the combination of the velocity law and density/optical depth profile described in Sect. 5.1. Thus, at any point in the 3-D space, we can infer the density and the velocity of the gas from our analytic framework. Here, we present and discuss the implications of the constraints on the mass outflow rate inferred from our outflow model.
The density of the outflowing gas is directly proportional to the inferred optical depth through Eq. 6. Replacing the constants, we have
$\tau_{0}=4.57\times 10^{-7}\mathrm{cm}^{3}\mathrm{s}^{-1}n_{0}\frac{R_{0}}{v_{0}},$ (25)
which leads to $n_{0}=5.76\times 10^{-6}$ cm-3, using the best-fitting model values of $\tau_{0}$, $R_{0}$, and $v_{0}$. However, in this context, $n_{0}$ corresponds to the number density of ionized Mg ($n_{\mathrm{Mg^{+}},0}$) at radius $R_{0}$. To compute the corresponding hydrogen density of the gas ($n_{\mathrm{H},0}$), it is necessary to correct by metallicity $\eta(\mathrm{Mg})$, dust depletion $d(\mathrm{Mg})$ and ionization fraction $\chi(\mathrm{Mg^{+}})$ (see, e.g., Martin et al., 2013). Specifically, $n_{\mathrm{H},0}$ is given by
$n_{\mathrm{H},0}=\frac{n_{\mathrm{Mg^{+}},0}}{\eta(\mathrm{Mg})\,\chi(\mathrm{Mg^{+}})\,d(\mathrm{Mg})}.$ (26)
@@
We used then the relation between $\chi(\mathrm{Mg^{+}})$ and $U$ reported by Martin et al. (2013), determined using Cloudy (Ferland et al., 2013), to compute the corresponding ionization fraction:
$\chi(\mathrm{Mg^{+}})=4\times 10^{-4}U^{-0.94},$ (27)
@@
$\dot{M}(r)=\mu\,m_{\mathrm{H}}\,n_{\mathrm{H}}(r)\,A(r)\,v(r),$ (28)
with $\mu\,m_{\mathrm{H}}$ equal to the average mass per hydrogen atom. We adopt a value of $\mu=1.4$, which is commonly used in this context (see, e.g., Rupke et al., 2002; Weiner et al., 2009; Martin et al., 2012; Burchett et al., 2021). Since we have imposed mass conservation, the expression above does not depend on the radius and can be written as
$\dot{M}=4\,\pi\,\mu\,m_{\mathrm{H}}\,n_{\mathrm{H},0}\,C_{\Omega}\,R_{0}^{2}\,v_{0},$ (29)
where $C_{\Omega}=1-cos(\mathrm{O.A.})$ corresponds to the angular covering fraction of the bicone. Evaluating the equation yields a mass outflow rate of
$\dot{M}\approx 12\pm 7\,\mathrm{M}_{\odot}\,\mathrm{yr}^{-1},$ (30)
@@
Overall, the measured outflow rate and mass loading factor could be up to a factor $\sim 6$ lower, if we assume a gas-phase metallicity given by the relation reported by Zahid et al. (2014), and boosted according to the findings from Chisholm et al. (2018) (with the caveat that a different metallicity calibration could potentially shift the relation from Zahid et al. (2014)). On the contrary, it could also be up to a factor $\sim 8$ higher, using the dust depletion reported by Savage & Sembach (1996). Despite the fact that we do not account for additional systematic variations induced by our ionization correction choice, the metallicity and dust depletion choices come with significant systematic uncertainties that must be taken into account in order to properly interpret our results and compare them with the results from different works.
@@
In the following, we proceed similarly to Burchett et al. (2021) to investigate whether this outflowing material is able to escape the galactic potential or will fall again into the galaxy. We use the stellar-to-halo mass relation from Moster et al. (2013) to calculate a total halo mass of $\log M_{\mathrm{h}}/\mathrm{M}_{\odot}=11.9$. For a halo of this mass with a Navarro–Frenk–White profile, and assuming a concentration parameter given by the relation from Ragagnin et al. (2019), we calculate an escape velocity of $\sim 250$ km s-1 at the maximum radius reached by the outflow of $\sim 70$ kpc. This is significantly lower than the terminal velocity of the outflow $v_{\mathrm{max}}=450$ km s-1, meaning that the outflowing material will likely escape the gravitational potential of the halo and enrich the intergalactic medium, rather than falling again into #884. However, as shown in previous sections, a radius of $70$ kpc is well beyond the radius where our outflow model describes the data faithfully. If we compare the escape velocity with the outflow velocity at a smaller radius, where our outflow model still performs well, we find that, for instance, at $25$ kpc, the escape velocity is still higher (although comparable, within the error bars) than the outflow velocity ($\sim 250$ and $\sim 220\pm 70$ km s-1, respectively). Specifically, given our set of best-fitting parameters, we find that the outflow starts exceeding the escape velocity of the galactic potential at $\sim 30$ kpc. Thus, whether the outflowing gas escapes the galactic potential or not depends on the radii up to which our outflow model provides an accurate representation of the data. Our analysis suggests that this holds for impact parameters of $\sim 20-25$ kpc. However, this cannot be translated directly into radii due to the additional line-of-sight component of the distance, and the face-on nature of the outflow. Moreover, we find evidence of high-velocity gas ($\sim 400$ km s-1) in the down-the-barrel spectrum of the outflow. Altogether, we find it more likely that our outflow model still performs well by the point where the outflow velocity overcomes the escape velocity of the galaxy halo and, thus, that the outflowing gas will be able to escape the galaxy.